# Capacity Gains in MIMO Systems with Few-Bit ADCs Using Nonlinear Analog Circuits

### v2

@@ II-B Summary of Prior Works
Prior Works on Analog Beamforming: Analog beamforming utilizes analog phase shifters and only one RF chain for the beamforming operation [16, 45]. This leads to a simplified design and low power consumption compared to hybrid and digital beamforming. However, analog beamforming can only support single-stream transmission which yields lower data rates.
@@ III Curse of Low Dimensions
Figure 4: Constellations resulting from linear processing prior to quantization in a scenario with $n_{t}=n_{r}=2$ and $n_{q}=4$.
Figure 5: Constellations resulting from nonlinear processing prior to quantization in a scenario with $n_{t}=n_{r}=2$ and $n_{q}=4$.
As mentioned in Section II, in both the digital and hybrid beamforming scenarios studied in prior works, there is a rate-loss associated with the use of low-resolution ADCs even in the high SNR regime. That is, while ideally one would wish to achieve a communication rate of $n_{q}$ bits per channel-use in the high SNR regime, in reality, the maximum achievable rate only grows logarithmically in the number of ADCs in both scenarios. In this section, we provide a high-level reasoning for this rate-loss. This intuitive explanation forms the motivation for the rest of the paper, and justifies the gains observed when using non-linear operators instead of linear operators as in hybrid beamforming schemes.
To explain the ‘curse of low dimensions’ phenomenon, let us consider a PtP MIMO system with hybrid beamforming at the receiver, operating in the high SNR regime, and consisting of $n_{t}$ transmit antennas, $n_{r}$ receive antennas, and $n_{q}$ one-bit ADCs. Given a hybrid beamforming matrix $\mathbf{V}$ and ADC threshold vector $\mathbf{t}$, one can define the quantization mapping $Q(\mathbf{Y})\triangleq\mathbf{\widehat{W}}=(sign(W_{1}),sign(W_{2}),\cdots,sign(W_{n_{q}}))$, where $\mathbf{W}=\mathbf{V}+\mathbf{t}$. An important characteristic of $Q(\cdot)$ is the resulting Voronoi regions — where a Voronoi region $\mathcal{P}$ is defined as the set of all vectors $\mathbf{Y}$ which are mapped to the same quantization output. We let $\mathsf{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\cdots,\mathcal{P}_{|\mathsf{P}|}\}$ denote the set of all Voronoi regions of $Q(\cdot)$. In the high SNR regime, it is straightforward to argue that the number of messages which can be transmitted reliably per channel-use is equal to the number of Voronoi regions of $Q(\cdot)$, i.e., the maximum achievable rate is $\log{|\mathsf{P}|}$ bits/channel-use. Ideally, the number of Voronoi regions grows exponentially with the number of one-bit ADCs (i.e., $|\mathsf{P}|=2^{n_{q}}$) so that rate equal to $n_{q}$ bits/channel-use is achieved. However, this is not the case for low-dimensional spaces where the maximum number of possible Voronoi regions produced by a mapping $Q(\cdot)$ may be significantly lower than $2^{n_{q}}$ [46]. In general, the maximum number of Voronoi regions is equal to the maximum number of regions in an $n_{r}$-dimensional space when cut by $n_{q}$ hyperplanes [47, 48] given by:
$|\mathsf{P}|=2\sum_{i=0}^{n_{r}-1}{n_{q}-1\choose i}.$ (1)
@@
Consequently, for a fixed $n_{r}$, the maximum achievable rate $\log{|\mathsf{P}|}$ given with (1) grows logarithmically with $n_{q}$. Whereas if $n_{r}=n_{q}$ (digital beamforming), then $\log{|\mathsf{P}|}$ grows linearly in $n_{q}$. We call this phenomenon the curse of low dimensions since it leads to reduced rates when the channel output dimension is low, i.e., when $n_{r}<n_{q}$. This is further clarified through the following example.
Example 1 (Curse of Low Dimensions).
@@
To provide a concrete example, let us consider a scenario with two transmit and receive antennas (i.e. $n_{t}=n_{r}=2$) and $n_{q}=4$ one-bit ADCs at the receiver. For ease of explanation, we focus on the achievable communication rates in the high SNR regime. In this case, the received signal space is a two-dimensional space consisting of points $(Y_{1},Y_{2})$. In hybrid beamforming, the received signals are passed through a linear analog combiner characterized by $\mathbf{V}=[v_{i,j}]_{i,j\in[4]}$, and the ADC thresholds are $t_{i}\in\mathbb{R},i\in[4]$. The ADCs perform the following comparisons $v_{i,1}Y1+v_{i,2}Y_{2}\lessgtr t_{i},i\in[4]$. The resulting Voronoi regions can be represented by a partition of the two-dimensional Euclidean space using four lines. For instance, in Phase-shift keying (PSK) modulation the thresholds are $t_{i}=0,i\in[4]$ and
@@
The resulting Voronoi regions are shown in Figure 4(a). It can be observed that there are 8 Voronoi regions, as opposed to the ideal 16 regions. Alternatively, Figure 4(b) shows the Quadrature Amplitude Modulation (QAM) Voronoi regions, and Figure 4(c) represents a general position constellation proposed in [30], which achieves the highest number of Voronoi regions possible using analog linear processing. As was observed in [30], and shown in the figure, hybrid beamforming architectures can achieve at most 11 unique Voronoi regions thus achieving a rate of $\log{11}$ bits/channel-use.
An alternative approach is to use non-linear analog processing instead of the linear processing used in hybrid beamforming architectures. For instance, absolute values of the signals can be generated in the analog domain using envelope detectors, and then using the ADCs, the following comparisons can be made:
$|\mathbf{Y}_{1}|\lessgtr 1,\quad|\mathbf{Y}_{2}|\lessgtr 1,\quad|\mathbf{Y}_{1}|\lessgtr 3,\quad|\mathbf{Y}_{2}|\lessgtr 3.$
The resulting Voronoi regions from such processing are shown in Figure 5(a). The regions whose elements are associated with unique ADC binary output vectors are indicated with a black circular marker. Note that not all regions result in a unique ADC output. For example, if $Y_{i}$ are very large positive numbers (top-right corner) or very large negative numbers (bottom-left corner), then the ADC outputs are $(1,1,1,1)$ for elements of both regions, hence only one of the two regions contributes to the achievable rate as it corresponds to a unique transmitted message. Similarly, the top-left and bottom-right regions produce ADC output $(1,1,1,1)$. In the figure, we have marked the top-left region with a black circular marker, which indicates its contribution to the rate computation. Thus, the rate is not defined by the total number of quantization regions but by the number of unique regions. The constellation produces a total of 9 unique quantization regions, resulting in a rate of $\log 9$ bits/channel use. Figure 5(b) shows a different set of Voronoi regions resulting from the following operations:
$|\mathbf{Y}_{1}|\lessgtr 1,\quad|\mathbf{Y}_{2}|\lessgtr 1,\quad|\mathbf{Y}_{1}+\mathbf{Y_{2}}|\lessgtr 1,\quad|\mathbf{Y}_{1}-\mathbf{Y}_{2}|\lessgtr 1,$
resulting in a rate of $\log 12$ bits/channel use and the constellation in Figure 5(c) results from the operations:
$|\mathbf{Y}_{1}|\lessgtr 1,\quad|\mathbf{Y}_{2}|\lessgtr 1,\quad|\mathbf{Y}_{1}-1|\lessgtr 1,\quad|\mathbf{Y}_{2}-1|\lessgtr 1,$
which results in the optimal rate of 4 bits/channel use, thus completely lifting the curse of low dimensions.
@@ IV SISO Systems Equipped with Envelope Detectors
As discussed in the introduction, two different approaches, namely hybrid beamforming and low resolution quantization, have been proposed to address the ADC power consumption in wideband systems. In this section, as a first step, we consider single-input single-output (SISO) systems — modeling analog beamforming — equipped with low resolution ADCs and evaluate the achievable rate using non-linear analog processing, particularly, envelope detectors. This serves as a foundation for the derivations for hybrid beamforming in the subsequent sections. Hence, we consider a MIMO receiver with analog beamforming equipped with a collection of envelope detectors for analog signal processing and a power budget of $P_{ADC}$ for ADC power consumption. Our main objective in this section is to characterize the optimal values for the number and resolution of the ADCs used at the receiver terminal, which achieve the maximum system throughput, given the signal-to-noise ratio of the channel and the ADC power budget.
IV-A Envelope Detectors for Non-linear Analog Processing
Envelope detectors are suitable for analog processing of signals at high frequencies due to their low power consumption and simple circuit design [49]. The power consumption of envelope detectors is significantly lower at high data rates (see Section VIII). An envelope detector is parametrized by its threshold $a\in\mathbb{R}$ and its operations on an input $x\in\mathbb{R}$ are captured by the function $A_{1}(x,a)=|x-a|,x\in\mathbb{R}$. Envelope detectors can be concatenated in a sequence to generate a larger collection of analog operators. The operation of a sequence of $s\in\mathbb{N}$ envelope detectors with bias vector $a^{s}=(a_{1},a_{2},\cdots,a_{s})$ is captured by the iterative relation $A_{s}(x,a^{s})=|A_{s-1}(x,a^{s-1})-a_{s}|,s>1$. It should be noted that concatenating large numbers of envelope detectors leads to increased circuit noise, and power consumption. Hence, there is a tradeoff between power-consumption, circuit complexity and robustness, and degrees of freedom in generating analog processing functions which in turn affects the set of achievable rates. This tradeoff is quantified in this section by deriving the set of achievable rates in terms of the number of concatenated envelope detectors.
IV-B Communication Strategies and Achievable Rates
In this section, we determine the optimal number and resolution of ADCs required at the receiver terminal to maximize system throughput, taking into consideration the signal-to-noise ratio of the channel and the ADC power budget. Additionally, we investigate the channel capacity for a given communication system, parametrized by $(P_{T},P_{ADC},\mathbf{h},\delta)$.

### v3

@@ II-C Conventional Beamforming Architectures
Analog Beamforming: Analog beamforming utilizes analog phase shifters and only one RF chain for the beamforming operation. This leads to a simplified design and low power consumption compared to hybrid and digital beamforming. However, it potentially yields lower rates. The analog processing function in analog beamforming is represented as:
$W=\mathbf{f}_{a}(\mathbf{Y})=\mathbf{b}^{T}\mathbf{Y},$
where $\mathbf{b}=(b_{1},b_{2},\cdots,b_{n_{r}})$ is the analog beamforming vector [^2] By default, we assume column vectors unless indicated otherwise. and $\mathbf{b}^{T}$ denotes the transpose of $\mathbf{b}$.
Hybrid Beamforming: Hybrid beamforming uses a collection of analog beamformers to linearly combine the received signals in the analog domain [30, 31]. This linear analog signal processing improves performance, in terms of achievable rates, by rotating the received signal such that the information-loss in the quantization step is reduced [32, 24]. The analog processing function in hybrid beamforming is represented as:
$\mathbf{W}=\mathbf{f}_{a}(\mathbf{Y})=\mathbf{v}\mathbf{Y},$ (2)
where $\mathbf{v}=[v_{i,j}]_{i\in[n_{q}],j\in[n_{r}]}$ is the hybrid beamforming matrix.
Digital Beamforming: In digital beamforming, each antenna is directly connected to its dedicated ADC, i.e., $n_{r}=n_{q}$ and $\mathbf{f}_{a}(\cdot)$ is the identity function [20, 33, 34]. This leads to an increased number of ADCs and potentially higher achievable rates at the expense of increased power consumption.
@@ III Curse of Low Dimensions
(a) PSK
(b) QAM
(c) General Position [27]
(d) Equation (5)
(e) Equation (6)
Figure 2: (a)–(c) The Voronoi Regions for various hybrid beamforming architectures considered in Section III-B. (d)-(e) Voronoi regions for beamforming architectures with nonlinear analog processing in the example of Section III-B. Blue-colored binary vectors correspond to regions with unique ADC outputs.
Figure 3: The Voronoi regions and ADC outputs produced when using envelope detectors in the example of Equation (4).
As discussed in the introduction, reducing the number and resolution of the ADCs leads to a reduction in the resulting channel capacity. This rate-loss is partially unavoidable as the limited number of output bits of the ADCs creates a fundamental bottleneck in the communication channel. For instance, if the receiver is equipped with a single one-bit ADC, then the communication rate cannot exceed one bit per channel-use even in the high SNR regime. More generally, for a receiver equipped with $n_{q}$ ADCs each with $\ell$ output levels, the communication rate is bounded from above by $n_{q}\log{\ell}$ bits per channel-use. This implies that the capacity may scale linearly with $n_{q}$. However, it has been observed in prior works that the channel capacity only scales logarithmically with $n_{q}$. For instance, in [20], it was shown that the high SNR capacity of the single-input multiple-output (SIMO) system with digital beamforming is well-approximated by $\log{(4n_{q}+1)}$. Similarly, the high SNR achievable rate in systems with hybrid beamforming grows logarithmically with $n_{q}$ [24]. In this section, we argue that this logarithmic growth, as opposed to the expected linear growth, is due to a phenomenon that we call the curse of low dimensions. We first provide an explanation of this phenomenon by evaluating the high-SNR capacity of hybrid beamforming systems equipped with one-bit ADCs. Then, we introduce a mitigating solution, using non-linear analog operators, and show its effectiveness through a numerical example. This intuitive explanation forms the motivation for the rest of the paper.
III-A High SNR Capacity of Hybrid Beamforming Systems
To provide a high-level description of the curse of low dimensions, let us consider the high SNR capacity of a hybrid beamforming MIMO system equipped with $n_{t}$ transmit antennas, $n_{r}$ receive antennas, $n_{q}$ one-bit ADCs, hybrid beamforming matrix $\mathbf{v}$, and ADC threshold vector $t_{i,1},i\in[n_{q}]$. Furthermore, let us assume that $\mathbf{Y}\approx\mathbf{h}\mathbf{X}$ in the high SNR regime. Hence, the channel output space is equal to the subspace generated by the columns of $\mathbf{h}$, denoted by $\mbox{Span}(\mathbf{h})$.
Each one-bit ADC yields a partition of the output space $\mbox{Span}(\mathbf{h})$ into two subspaces, where one subspace corresponds to points yielding an ADC output equal to zero, and the other corresponds to points yielding ADC output equal to one. To elaborate, recall from Section II-C that in hybrid beamforming the input to the $i$th ADC is given by $W_{i}=\mathbf{v}_{i,:}\mathbf{Y}$. Then, the $i$th ADC output is given by
$\widehat{W}_{i}=\begin{cases}0\qquad&\text{ if }W_{i}<t_{i,1},\\ 1&\text{otherwise}\end{cases}.$
We define the decision region $\mathcal{S}_{i,0}=\{\mathbf{y}|w_{i}\leq t_{i}\},i\in[n_{q}]$, and its complement as $\mathcal{S}_{i,1}=\mathcal{S}_{i,0}^{c}$, where $w_{i}=\mathbf{v}_{i,:}\mathbf{y}$. The output vector of the ADC module is the binary-valued vector $\mathbf{\widehat{W}}=(\widehat{W}_{1},\widehat{W}_{2},\cdots,\widehat{W}_{n_{q}})$, where $\mathbf{Y}\in\mathcal{S}_{i,\widehat{W}_{i}},i\in[n_{q}]$. Thus, for a given binary vector $\mathbf{\widehat{w}}\in\{0,1\}^{n_{q}}$, we have that
$\widehat{\mathbf{W}}=\widehat{\mathbf{w}}\iff\mathbf{Y}\in\bigcap_{i\in[n_{q}]}\mathcal{S}_{i,\widehat{w}_{i}}.$
Each non-empty $\bigcap_{i\in[n_{q}]}\mathcal{S}_{i,\widehat{w}_{i}},\mathbf{\widehat{w}}\in\{0,1\}^{n_{q}}$ is called a Voronoi region. Let $C_{h}$ be the high SNR capacity. Then, by the information processing inequality [35], we have that:
$C_{h}=\max_{P_{\mathbf{X}}}I(\mathbf{X};\mathbf{\widehat{\mathbf{W}}})\leq\log{|\mathcal{\widehat{W}}|},$
where $\mathcal{\widehat{W}}\!=\!\{\mathbf{\widehat{w}}|\bigcap_{i\in[n_{q}]}\!\mathcal{S}_{i,\widehat{w}_{i}}\!\neq\!\varnothing\}\!$ is the set of all Voronoi regions.
Ideally, we expect $|\mathcal{\widehat{W}}|=2^{n_{q}}$, so that rate is equal to $n_{q}$ bits per channel-use. However, this is not the case for low-dimensional spaces. To explain further, let the dimension of the subspace $\mbox{Span}(\mathbf{h})$ be denoted by $n_{d}$. Then, the maximum number of partition regions $|\mathcal{\widehat{W}}|$ is equal to the maximum number of regions in an $n_{d}$-dimensional space when cut by $n_{q}$ hyperplanes [36]:
$|\mathcal{\widehat{W}}|\leq 2\sum_{i=0}^{n_{d}-1}{n_{q}-1\choose i}.$ (3)
@@ III-A High SNR Capacity of Hybrid Beamforming Systems
Note that (${n_{q}-1\choose i}\leq(n_{q}-1)^{i}<n_{q}^{n_{d}}$ for $i\in\{0,1,\cdots,n_{d}-1\}$. Consequently, for a fixed $n_{d}$, the maximum achievable rate is $\log{|\mathcal{\widehat{W}}|}<n_{d}\log{n_{q}}$, and grows logarithmically with $n_{q}$. This is also separately observed in prior works [20, 24, 27]. On the other hand, if $n_{d}=n_{q}$, then $\log{|\mathcal{\widehat{W}}|}$ grows linearly with $n_{q}$. We call this phenomenon the curse of low dimensions since it leads to reduced rates when the channel output dimension is smaller than the number of ADCs, i.e., when $n_{d}<n_{q}$.
III-B A Mitigating Solution Using Non-Linear Processing
@@
In this work, we argue that non-linear analog processing prior to quantization, using envelope detectors and polynomial operators, mitigates the aforementioned rate-loss. To provide a concrete example, let us consider a high SNR communication scenario with $n_{t}=n_{r}=2$ and $n_{q}=4$ one-bit ADCs at the receiver. Further assume that the channel is full-rank, so that $n_{d}=2$. The ADCs perform the following comparisons $v_{i,1}Y_{1}+v_{i,2}Y_{2}\lessgtr t_{i},i\in[4]$. The Voronoi regions can be represented as a partition of the two-dimensional Euclidean space using four lines. For instance, in phase-shift keying (PSK) modulation the thresholds are set to $t_{i}=0,i\in[4]$ and the linear combining coefficients are:
@@
As shown in Figure 2(a), there are $|\mathcal{\widehat{W}}|=8$ Voronoi regions, as opposed to the ideal 16 regions. Alternatively, Figure 2(b) shows the quadrature amplitude modulation (QAM) Voronoi regions, and Figure 2(c) represents a general position constellation proposed in [27], which achieves the highest number of Voronoi regions possible using hybrid beamforming [27]. Consequently, PSK achieves a high-SNR rate of at most $\log{8}$, QAM achieves $\log{9}$, and the construction of [27] achieves $\log{11}$ bits per channel-use.
We propose to use non-linear analog processing to mitigate the aforementioned rate-loss. To elaborate, assume that the receiver can perform absolute value operations in the analog domain using envelope detectors. In particular, it receives $(Y_{1},Y_{2})$ and produces the analog signals $W_{1}=|Y_{1}|,W_{2}=|Y_{1}-1|,W_{3}=|Y_{2}|$ and $W_{4}=|Y_{2}-1|$. It then feeds each of these four signals to a one-bit ADC with threshold equal to one. That is, the ADC module performs the following comparisons:
$|Y_{1}|\lessgtr 1,\quad|Y_{1}-1|\lessgtr 1,\quad|Y_{2}|\lessgtr 1,\quad|Y_{2}-1|\lessgtr 1.$ (4)
The resulting Voronoi regions are shown in Figure 3. Note that the resulting partition is a five-by-five grid in the two-dimensional space, which yields a total of 25 partition elements. However, the ADC outputs are not unique for all of these partition elements. For instance, the top-left partition element, which corresponds to $Y_{1}\to-\infty$ and $Y_{2}\to\infty$, produces ADC output sequence $\widehat{\mathbf{W}}=(0,0,0,0)$. Similarly, the bottom-right partition element, which corresponds to $Y_{1}\to\infty$ and $Y_{2}\to-\infty$, yields ADC output sequence $\widehat{\mathbf{W}}=(0,0,0,0)$. As shown in the figure, there are $|\widehat{\mathcal{W}}|=16$ partition elements with unique ADC outputs, yielding an (optimal) high SNR rate of $\log{16}=4$ bit per channel-use.
To provide further insights into how the choice of non-linear analog processing functions affects the resulting channel capacity, we have provided two additional examples in Figure 2. In particular, Figure 2(d) shows the Voronoi regions using zero-bias envelope detectors:
$|{Y}_{1}|\lessgtr 1,\quad|{Y}_{2}|\lessgtr 1,\quad|{Y}_{1}|\lessgtr 3,\quad|{Y}_{2}|\lessgtr 3,$ (5)
and Figure 2(e) shows the Voronoi regions resulting from using analog linear combining and zero-bias envelope detectors:
$|{Y}_{1}|\lessgtr 1,\quad|{Y}_{2}|\lessgtr 1,\quad|{Y}_{1}+{Y_{2}}|\lessgtr 1,\quad|{Y}_{1}-{Y}_{2}|\lessgtr 1.$ (6)
It can be observed that the constellation of Figure 2(e) achieves a high SNR capacity of $\log{9}$ bits per channel-use, whereas that of 2(d) achieves $\log{12}$ bits per channel-use.
@@ IV SISO Systems Equipped with Envelope Detectors
In this section, we consider a single-input single-output (SISO) system, where the analog processing operators are restricted to concatenated sequences of envelope detectors. The (scalar) channel input and output are related through $Y=hX+N$, where $N\sim\mathcal{N}(0,1)$. We denote the transmit power by $P_{T}$ and the ADC power budget by $P_{ADC}$. The maximum number of concatenated envelope detectors is denoted by $\delta$. We represent such a system by the tuple $(P_{T},P_{ADC},{h},\delta)$. The resulting channel capacity is represented by $C_{env}(P_{T},P_{ADC},{h},\delta)$.
IV-A Concatenated Sequences of Envelope Detectors
We model the operation of an envelope detector as an absolute value operation. For instance, an envelope detector with input $y\in\mathbb{R}$ and bias $b\in\mathbb{R}$ outputs $|y-b|$. Sequences of envelope detectors can be concatenated with each other. For instance, consider a concatenated pair of envelope detectors with biases $b_{1},b_{2}\in\mathbb{R}$, respectively. Then, for input $y\in\mathbb{R}$, the output of the concatenated pair is $\big{|}|y-b_{1}|-b_{2}\big{|}$. Due to analog circuit limitations, we assume that a maximum of $\delta\in\mathbb{N}$ envelope detectors can be concatenated. Consequently, we define the set of implementable analog functions as:
$\mathcal{F}^{\delta}_{env}=\{f(y)=A_{s}(y,b^{s}),y\in\mathbb{R}|s\in[\delta],b^{s}\in\mathbb{R}^{s}\},\quad\delta\in\mathbb{N},$
where $A_{1}(y,b)\triangleq|y-b|,y,b\in\mathbb{R}$ and $A_{s}(y,b^{s})\triangleq A_{1}(A_{s-1}(y,b^{s-1}),b_{s})=|A_{s-1}(y,b^{s-1})-b_{s}|,s\in\mathbb{N}$. That is, $\mathcal{F}^{\delta}_{env}$ consists of all functions generated using sequences of $s\leq\delta$ concatenated envelope detectors. Implementability is discussed in Section VIII. Concatenating large numbers of envelope detectors leads to increased circuit noise and power consumption. Hence, there is a tradeoff between power-consumption, circuit complexity and robustness, and degrees of freedom in implementing analog processing functions.
IV-B The Quantizer and Its Associated Codebook
We call the combination of the envelope detectors and the ADC module the quantizer, and represent its operation by a function $\mathbf{q}:\mathbb{R}\to[\ell]^{n_{q}}$, where $n_{q}$ is the number of ADCs and $\ell$ is the number of output levels of each ADC, i.e., $\mathbf{q}(y)=\mathbf{a}(\mathbf{f}_{a}(y))$. This is clarified through the following example.
Example 1 (Quantizer).
Consider a SISO system equipped with two ADCs with three output levels, i.e., $n_{q}=2$ and $\ell=3$. Assume that a maximum of two envelope detectors can be concatenated with each other, i.e., $\delta=2$. Let $f_{1}(y)=A_{2}(y,(2,3))=||y-2|-3|$ be the input to the first ADC and $f_{2}(y)=A_{2}(y,(0,4))=||y|-4|$ the input to the second ADC. Let the ADC thresholds be given by:
$t_{1,1}=1,\quad t_{1,2}=2,\quad t_{2,1}=1,\quad t_{2,2}=2,$
The quantization function is $\mathbf{q}(y)=(\widehat{w}_{1},\widehat{w}_{2})$, where $\widehat{w}_{i},i\in\{1,2\}$ are the two ADC outputs. For instance, for $y=-1$ the ADC inputs are $f_{1}(-1)=0$ and $f_{2}(-1)=3$. Since $f_{1}(-1)<t_{1,1}$ and $f_{2}(-1)>t_{2,2}$, the ADC output vector is $\mathbf{q}(-1)=(0,2)$. For general $y\in\mathbb{R}$, the quantizer output $\mathbf{q}(y)$ is shown in the last row of the table in Figure 4.